### v1

# Deep Research Comparator: A Platform For Fine-grained Human Annotations of Deep Research Agents

Abstract
Effectively evaluating deep research agents that autonomously search the web, analyze information, and generate reports remains a major challenge, particularly when it comes to assessing long reports and giving detailed feedback on their intermediate steps. To address these gaps, we introduce Deep Research Comparator, a platform that offers a holistic framework for deep research agent hosting, side-by-side comparison, fine-grained human feedback collection, and ranking calculation. Given a user query, our platform displays the final reports from two different agents along with their intermediate steps during generation. Annotators can evaluate the overall quality of final reports based on side-by-side comparison, and also provide detailed feedback separately by assessing intermediate steps or specific text spans within the final report. Furthermore, we develop Simple Deepresearch, an end-to-end agent scaffold. This scaffold serves as a baseline that facilitates the easy integration of various large language models to transform them into deep research agents for evaluation. To demonstrate the platform’s utility for deep research agent development, we have collected real user preference data from 17 annotators on three deep research agents. We plan to release the code base and collected data.  [^1] Our code is available at GitHub, and demo video can be found at Youtube.
Deep Research Comparator: A Platform For Fine-grained Human Annotations of Deep Research Agents
Prahaladh Chandrahasan1∗  Jiahe Jin1∗  Zhihan Zhang1∗ Tevin Wang1Andy Tang1Lucy Mo1Morteza Ziyadi2Leonardo F.R. Ribeiro2 Zimeng Qiu2Markus Dreyer2Akari Asai1Chenyan Xiong1 1Carnegie Mellon University  2Amazon {prahalac,jjiahe,zhihanz,tevinw,andyt,linmo}@andrew.cmu.edu {mziyadi,leonribe,zimengqi,mddreyer}@amazon.com aasai@cs.cmu.educx@cs.cmu.edu
11footnotetext: Equal Contribution.
1 Introduction
Recent advancements in large language models (LLMs) have catalyzed the development of autonomous agents, which can utilize various tools to solve real-world tasks. Among them, the deep research agent has recently emerged, referring to an agent powered by LLMs that autonomously finds, analyzes, and synthesizes vast amounts of online information to produce a comprehensive report Xu and Peng (2025). Prominent proprietary examples include agents developed by OpenAI OpenAI (2025), Google Gemini Gemini (2025), and Perplexity perplexity (2025), and the open-source community has also produced notable contributions Li et al. (2025); GPTResearcher (2025).
Figure 1: The main interface of Deep Research Comparator. Users can submit their query in the text box, and the intermediate steps and final reports from two agents will be displayed side by side.
Despite these advancements, evaluating the long-form reports generated by deep research agents remains a significant challenge. Existing benchmarks for search agents Wei et al. (2025); Wu et al. (2025) primarily target short-form answers, leaving the assessment of long-form reports largely underexplored. Moreover, unlike tasks such as math or coding, there are no widely accepted objective metrics for assessing long reports (Asai et al., 2024). To this end, some existing work focuses on developing hand-crafted rubrics for evaluation Gou et al. (2025); FutureSearch et al. (2025); Arora et al. (2025), but these approaches are difficult to scale.
Beyond the difficulty in assessing long reports, we also lack a fine-grained evaluation of agents’ intermediate steps. LLM-based agents often execute multi-step actions to complete long-horizon tasks, but current evaluations are typically based on outcomes Deng et al. (2023); Zhou et al. (2024); Jimenez et al. (2024); Xie et al. (2024), resulting in sparse and delayed feedback. However, process feedback is crucial for training methods like reinforcement learning (RL) to accurately attribute step contribution in the trajectory Feng et al. (2025); Wang et al. (2025); Setlur et al. (2025). Such feedback is also essential for understanding agent behavior and assessing the reliability of the generated reports Lightman et al. (2023); Wan et al. (2024).
To address these gaps, we introduce Deep Research Comparator (Section 3), a platform for fine-grained human evaluation of long-report generation by deep research agents. It supports easy integration of new agents with minimal engineering effort, enabling both open-source and closed-source systems to be hosted through a unified web interface. When a user submits a query, the platform randomly routes it to two agents and presents their intermediate steps and final reports.
Deep Research Comparator enables the collection of two types of human feedback. First, users can perform a side-by-side comparison of the two final reports and vote for the better one (Figure 1 bottom), similar to Chatbot Arena Chiang et al. (2024). These preference votes can then be used to calculate an outcome-based ranking of the agents. Second, users can provide more fine-grained annotations by upvoting or downvoting individual intermediate steps and specific text spans within the reports to indicate their quality (Figure 2). This detailed feedback creates a valuable dataset for potential future use, such as agent fine-tuning or reward modeling.
To support easy deployment and evaluation of different LLMs as deep research agents, we developed Simple Deepresearch, an end-to-end agent scaffold serving as a baseline (Section 4). This prompt-based scaffold allows various LLMs to be easily integrated, transforming them into deep research agents for evaluation.
Using our platform, we conducted a proof-of-concept experiment on three agents with 176 user queries (Section 5). We then ranked the agents based on pairwise comparison votes on their final reports. Additionally, we collected 1,281 upvote/downvote feedback on intermediate steps and 593 upvote/downvote with highlighted text spans in reports. From this, we calculated each agent’s overall upvote rate on intermediate steps for a process-based evaluation metric. Based on the collected data, we derive some design considerations for future agent developers on how to effectively use our platform for annotation and evaluation.
Overall, we make three key contributions:
  1. 1.
    We provide Deep Research Comparator, an open source platform to host deep research agents and supports fine-grained human evaluation through a unified web interface.
  2. 2.
    We develop Simple Deepresearch, a baseline agent scaffold that allows for the easy integration of various LLMs into deep research agents for evaluation.
  3. 3.
    We demonstrate the utility of our platform by collecting 176 real-world user queries with comparison votes, and fine-grained annotation on 1,281 steps and 593 text spans in reports.
2 Related Work
Figure 2: Demonstration of fine-grained annotation. Left: Users can upvote/downvote each step in the generation process. Right: Users can highlight any text span in the final report and upvote/downvote it.
LLM-based search agents
Search Agents are LLM-based systems that autonomously and iteratively use search tools to gather external information Jin et al. (2025); Schick et al. (2023); Anthropic (2025); Chen et al. (2024) for solving complex search tasks. These agents can generate dynamic action decisions that are executed as real-time internet searches, and then synthesize this information to provide up-to-date and relevant responses. Among them, “deep research agents” represent a more capable evolution of this paradigm OpenAI (2025); GPTResearcher (2025); Li et al. (2025), which conduct substantially more searches, engage in long-horizon planning, and operate with significantly higher test-time compute budgets, for more demanding, research-intensive scenarios.
Evaluations for Search Agents
Evaluations for search agents can be roughly divided into two types: short-answer tasks and long-form report generation. Short-answer tasks are easier to evaluate due to the availability of definitive ground truth. Commonly employed evaluations include intricate tasks that demand extensive knowledge and complex reasoning, such as GPQA Rein et al. (2023) and Human’s Last Exam HLETeam (2025). Specialized benchmarks for web searches have also been developed Mialon et al. (2023); Wei et al. (2025); FutureSearch et al. (2025); Wu et al. (2025), with BrowseComp Wei et al. (2025) as a recent representative. Compared to short-answer tasks, evaluating long-form reports presents a greater challenge. Some efforts have introduced hand-crafted, rubric-based benchmarks Gou et al. (2025); Arora et al. (2025); FutureSearch et al. (2025); Asai et al. (2024). Other studies have explored using LLMs to automatically score long reports against predefined criteria Coelho et al. (2025); Li et al. (2025).
3 Deep Research Comparator
This section presents Deep Research Comparator, a platform for fine-grained human evaluation of diverse deep research agents.
3.1 User Interface and Evaluation Workflow
Our platform provides an interactive, web-based user interface designed for side-by-side, fine-grained human annotation, as shown in Figure 1.
Annotators can submit arbitrary queries, and our system will randomly assign the query to two anonymous deep research agents. As the report generation begins, the agents’ intermediate steps are streamed to the front-end, followed by the final reports. After both reports are generated, users vote for their preferred report by selecting one of four choices: Agent A, Agent B, Tie, or Both are bad. These pairwise comparisons are ultimately used to compute the outcome-based rankings.
Besides evaluating the overall quality of the final reports, this platform also supports more fine-grained evaluation (Figure 2). Each step of the agent is accompanied by upvote and downvote icons, allowing users to provide step-wise feedback. As a long report may contain content of varying quality, our platform also supports highlighting specific text spans in the final report and providing feedback either as upvotes or downvotes.
3.2 System Architecture
Figure 3: High-level overview of Deep Research Comparator’s system architecture.
Figure 3 presents a high-level overview of our architecture. Our platform supports hosting a wide range of deep research agents. Agents that stream both intermediate steps and a final report in accordance with a unified JSON-format interface (see Appendix B) can be integrated into the platform with minimal engineering effort. This design allows agent developers to flexibly customize the content of intermediate steps to accommodate their annotation requirements and imposes no constraints on the agent’s internal architecture.
Frontend
The frontend presents the generated intermediate steps and final reports in a consistent format on the user interface. It is implemented as a static web application, with static files produced during the build process and hosted in an AWS S3 bucket, served to users through CloudFront CDN caching ensuring low latency.
Main Backend Service
The system routes user queries from the frontend to the Main Backend Service hosted on the main CPU node. It is responsible for handling user requests, selecting agents, routing queries to the selected agents, processing user votes, and serving the intermediate steps, final reports, and rankings back to the frontend.
Agent Serving Service
User queries are routed to the Agent Serving Services deployed on dedicated agent CPU nodes, each encapsulated within a Docker container. These agent services stream their responses back to the main backend via internal API calls using the unified JSON format specified in Appendix B. For open-source agent services whose underlying LLMs require GPU resources, the LLMs are deployed on specialized GPU nodes and accessed by the agent services through internal API calls. For closed-source agents, specialized proxy pods are implemented to invoke the agents through external APIs.
Ranking Calculation Job
This CronJob reads agent conversations and user annotation data from the database, and computes both outcome-based evaluation metrics (such as model rankings and query counts) and process-based metrics (such as step upvote rate). It then writes the computed results back to the relational database, which are served to the user interface via the Main Backend Service.
4 Simple Deepresearch
Simple Deepresearch (Figure  4) is a prompt-based, end-to-end agent scaffold that allows for the easy integration of different LLMs via the OpenAI API, transforming it into a deep research agent. It serves as a baseline for evaluation.
Figure 4: Overview of Simple Deepresearch, our Baseline agent scaffold.
The agent’s workflow is an iterative process. At step $k$, the agent uses the user query $q$ and the history $h_{k}$ as input, and generates the thought process $t_{k}$ and action $a_{k}$. Each action consists of an action type and corresponding content, adhering to a predefined action space (see Table 1) specified in the agent’s system prompt (see Appendix A). If $a_{k}$ is a search action, the agent queries a search API and receives documents as observation $obs_{k}$. Other actions do not involve interaction with the external environment but instead serve to update the agent’s memory by being incorporated into the history of next step.
Table 1: The action space of Simple Deepresearch.
Type Content
plan The plan for the research process.
search Web search query.
script Draft for the final report.
summary Summary of history steps.
answer The final report.
The history $h_{k}$ contains the information of the preceding steps (from 1 to $k-1$), providing the agent with the necessary context for its current step. If $a_{k}$ is an answer action, then the generation process ends. Otherwise, the history for step $k+1$ is updated as follows:
$h_{k+1}=\begin{cases}h_{k}+t_{k}+a_{k}+obs_{k},&a_{k}=\texttt{Search}\\ h_{k}+t_{k}+a_{k},&a_{k}=\texttt{Plan, Script}\\ a_{k},&a_{k}=\texttt{Summary}\end{cases}$
The summary action is introduced to better accommodate models with shorter context lengths. This allows the model to summarize the entire history, preserving the information it considers most important. The action content of summary action will be used as history for the next step.
5 Experiments
5.1 Human Annotation
We conduct human annotation experiments on Deep Research Comparator as proof-of-concept, with experimental setup and evaluation details provided below.
Agents
We evaluated three deep research agents: (1) the proprietary system Perplexity Deep Research perplexity (2025), (2) the open-source  GPT Researcher GPTResearcher (2025), configured in deep research mode with GPT-4.1 as the smart LLM, o4-mini as the strategic LLM, and Serper as search API (3) Simple Deepresearch(Gemini 2.5 Flash), our baseline agent scaffold intergrated with Gemini 2.5 Flash Deepmind (2025) as the underlying LLM and uses ClueWeb22 Overwijk et al. (2022) as search API.
Annotators
A total of 17 university students and industry researchers participated in the data annotation process. To measure inter-annotator consistency, we also provided annotators with four fixed queries (Appendix C), which allowed us to compute human agreement on a controlled set.
Metrics
Following Chatbot Arena Chiang et al. (2024), we use the Bradley-Terry model Bradley and Terry (1952) to aggregate pairwise comparison votes and compute a ranking based on user preferences for final reports. The final Bradley-Terry (BT) scores are computed relative to our baseline- Simple Deepresearch, which is assigned a fixed score of 1000. In addition to the outcome-based BT score, we also introduce a process-based evaluation metric: the step upvote rate. For each agent, we aggregate all upvotes and downvotes received by its intermediate steps and compute the step upvote rate as $\text{upvotes}/(\text{upvotes}+\text{downvotes})$.
5.2 Evaluation on Static Benchmark
To validate the performance of our baseline agent scaffold and compare human evaluation results with those from static benchmarks, we evaluated agents on DeepResearchGym Coelho et al. (2025), a benchmark that uses LLM-as-a-judge to assess long reports generated by deep research agents.
Our evaluation employs two key metrics targeting different aspects of writing quality and analytical depth. We provide the LLM judge with detailed evaluation criteria and prompt it to assign scores based on it. Specifically, Clarity assesses the logical coherence and linguistic fluency of the report, while Insight captures the analytical nuance and depth of reasoning demonstrated.
5.3 Results
The annotation experiment ultimately yields 176 queries across a diverse range of domains—including science, technology, media, society, economics, health, and culture—and are complex enough to require extensive web searches to answer (See Appendix D for an example). The win rate matrix between each pair of agents is calculated based on comparison votes on these queries, as shown in Figure 5. The Bradley-Terry (BT) scores and the final ranking derived from them are presented in the first two columns of Table 2.
In addition to the side-by-side comparison votes, we collect 1,281 instances of upvote/downvote feedback on intermediate steps, and 593 instances with highlighted text spans within the reports. The final step upvote rate is shown in the third column of Table 2, and detailed statistics for fine-grained feedback are presented in Table 3.
Table 2: Evaluation results of deep research agents: human evaluation on Deep Research Comparator(Ranking, BT Score, Step Upvote Rate) and LLM-based evaluation on DeepResearchGym (Clarity, Insight).
Agent Ranking BT Score Step Upvote Rate (%) Clarity Insight
GPT Researcher 1 1135.28 74.66 89.8 90.4
Perplexity DeepResearch 2 1087.41 90.72 89.5 89.3
Simple Deepresearch (Gemini 2.5 Flash) 3 1000.00 90.33 89.4 79.4
Table 3: Upvote/downvote counts for fine-grained annotation. PD: Perplexity DeepResearch, GR: GPT Researcher, SD: Simple Deepresearch (Gemini 2.5 Flash).
Intermediate Step Text Span
Agent Upvote Downvote Upvote Downvote
GR 255 33 445 151
PR 161 17 352 36
SD 101 27 271 29
Figure 5: Win Probability of Agent A over Agent B.
The results of static benchmark evaluation for agents involved in the human annotation are presented in the last two columns of Table 2, and results for additional agents are in Table 3. Results for human agreement are presented in Table 4.
5.4 Discussion
Step-Level Feedback Analysis
Our system displays agents’ intermediate steps, with their content determined solely by agent designers. As a result, the content and granularity of steps vary across agents. Notably, although GPT Researcher ranked highest in outcome-based evaluation, it received the lowest step—upvote rate, likely due to its multi-agent parallel framework, which produces loosely ordered and often abstract or minimal step messages. In contrast, Simple Deepresearch (Gemini 2.5 Flash) present detailed steps that include both thought and actions (Appendix 2. Intermediate Steps), contributing to a higher upvote rate. This highlights the importance for agent designers to consider what step content to expose, tailoring the level of detail and structure to their annotation goals.
LLM and Human Evaluation Agreement
For report quality assessment, human rankings closely align with the LLM-based evaluations on Clarity and Insight. However, human agreement was relatively low. This may be due to the small number of queries and the fact that all three agents perform at comparably high levels (see Table 6), making fine distinctions difficult. This suggests two directions for annotation. First, as LLMs grow more capable, advancing agent performance in specialized domains may require expert annotations to provide high-quality guidance. Second, the subtlety of human preferences highlights the value of broader community annotations when the goal is to align agent behavior with general user expectations.
6 Conclusion
We present Deep Research Comparator, an open-source platform that offers a holistic framework for deep research agent hosting, web front-end demonstration, side-by-side evaluation, fine-grained annotation, and also ranking calculation. We also introduce Simple Deepresearch, a baseline scaffold that allows the easy integration of different LLMs to evaluate them as deep research agents.
Using our platform, we collected votes for 176 user queries and computed rankings based on them, and also gathered fine-grained annotation for 1,281 intermediate steps and 593 text spans in reports. Our platform can benefit a wide range of downstream applications, such as benchmarking, agent behavior analysis, and agent training via techniques like process supervision and RLHF. We believe Deep Research Comparator will serve as a useful resource for the deep research agent community.
Acknowledgements
We would like to extend our thanks Amazon AGI for providing us with the required infrastructure support for hosting Deep Research Comparator.
Appendix A Simple Deepresearch Scaffold Prompt
You are a research assistant with the ability to perform web searches to write a comprehensive scientific research article in markdown format. You will be given a question, and you will need to write a report on the question. You can use search tools to find relevant information.
You dont need to write the report in one turn. You can search and revise your report multiple times. When you consider you need some new information, you can perform a search action. When you want to update, generate, or revise your report scripts, you can perform a scripts action. When you consider you have enough information, you can output the final report.
Based on the history information, you need to suggest the next action to complete the task.
You will be provided with:
1. Your history turns information: it might contains your previous plan, report scripts, search results. For search results, queries are in format <search> query </search> and the returned search results in <information> and </information>.
2. The question to answer.
IMPORTANT: You must strictly adhere to the following rules:
1. Choose ONLY ONE action from the list below for each response, DO NOT perform more than one action per step.
2. Follow the exact syntax format for the selected action, DO NOT create or use any actions other than those listed.
3. **Dont do duplicate search.** Pay attention to the history search results.
4. **Do not always perform the search action. You must consider the history search results and update your report scripts.**
Valid actions:
1. <search> query </search>: search the web for information if you consider you lack some knowledge.
2. <plan> plan </plan>: plan the report in your first turn.
3. <scripts> revised or newly generated report scripts </scripts>: revise former report scripts, or newly generate report scripts.
4. <summary> important parts of the history turns </summary>: summarize the history turns. Reflect the plan, scripts, search queries, and search results in you history turns, and keep the information you consider important for answering the question and generating your report. Still keep the tag structure, keep plan between <plan> and </plan>, keep scripts between <scripts> and </scripts>, keep search queries between <search> and </search>, and keep search results between <information> and </information>. The history turn information for your subsequent turns will be updated accoring to this summary action.
5. <answer> final report </answer>: output the final report.
Question: {question}
History Turns: (empty if this is the first turn)
Appendix B Unified JSON format
# Standardized fields that all agents must support:
"intermediate_steps": str,
"final_report": str,
"is_intermediate": bool,
"is_complete": bool,
"citations": List[str]
Appendix C Fixed Question Set
  1. 1.
    How have historical trends in Spain’s real estate market influenced the development and investment potential of golf properties?
  2. 2.
    How does the inclusion of personal life details in journalist biographies affect public perception of their professional credibility and integrity?
  3. 3.
    How does product lifecycle management affect sustainability practices within the remote car starter industry?
  4. 4.
    What role does digital personalization play in enhancing user engagement with their digital devices, particularly among anime fans?
Appendix D Annotation Example
This is a annotation example from our human annotation experiment:
Question: How does the inclusion of personal life details in journalist biographies affect public perception of their professional credibility and integrity?
Session ID: 24bed070-c202-46f3-ba23-37cf42b88f59
Agent A: Simple Deepresearch (Gemini 2.5 Flash)
Agent B: Perplexity DeepResearch
Final Vote: Agent B
Fine-grained Annotation
1. Text Spans
Agent: Perplexity DeepResearch
Text Span 1 “The inclusion of personal life details in journalist biographies represents a complex intersection of transparency efforts, branding strategies, and ethical considerations in contemporary journalism.” Vote: Upvote
Text Span 2 “Research indicates that while personal disclosures can enhance audience familiarity with reporters, they yield minimal impact on perceptions of professional credibility and integrity.” Vote: Upvote
2. Intermediate Steps
Agent: Simple Deepresearch (Gemini Flash 2.5)
Intermediate Step (Part 1) Step 1 Thought: **Planning a Deep Dive into Journalists’ Credibility** Okay, so I’m tackling this interesting question: How do the personal lives of journalists impact how the public perceives their professionalism? I’m already formulating my plan. First, I need to get organized. I’m thinking of structuring this research article in a way that builds a solid argument. My initial approach is to start with a strong **introduction**, setting the stage and clearly defining the scope of my investigation. Then, I’ll delve into a thorough **literature review**, laying out the existing studies and relevant theoretical frameworks that inform this complex relationship. Next, since this isn’t an original study, it will be a deep dive into an **analysis of existing methodology** and existing research methodologies. I need to focus on what research has already been completed, and what are its shortcomings. ACTION : Plan: 1. **Introduction:** * Define "journalist biographies" in this context (e.g., "About Me" pages, social media profiles, author bios in articles). * State the core question: How does the inclusion of personal life details affect public perception of journalist credibility and integrity? * Briefly outline the importance of the topic (e.g., trust in media, role of journalists in society). * Provide a roadmap for the article.
Intermediate Step (Part 2) 2. **Theoretical Framework/Literature Review:** * **Credibility:** Define and discuss established models of source credibility (e.g., expertise, trustworthiness, goodwill). * **Integrity:** Define journalistic integrity (e.g., objectivity, fairness, accuracy, ethical conduct). * **Self-Disclosure Theory:** Explore how self-disclosure can impact perceptions and relationship building. * **Parasocial Interaction Theory:** Discuss how audiences form relationships with media figures and how personal details might foster or hinder this. * **Impression Management Theory:** How journalists might use personal details to manage their public image. * Review existing studies (if any) that touch upon personal revelations by professionals (not just journalists) and public perception. Vote: Upvote
In our preliminary data annotation experiment, the interface only had two buttons for vote: “Agent A is better” and “Agent B is better”. But in the final platform we release, all four buttons are available (“Agent A is better”, “Agent B is better”, “Tie”, and “Both are bad”).
Table 4: Human agreement results on fixed queries, measured by consensus rate. All submissions on our platform are anonymous, as the system does not require user login. Note: PD = Perplexity DeepResearch, SD = Simple Deepresearch (Gemini 2.5 Flash), GR = GPT Researcher. Green indicates the majority-voted agent, while red indicates the minority.
Query GPT Researcher vs Perplexity DeepResearch Simple DeepResearch (Gemini 2.5 Flash) vs GPT Researcher Simple DeepResearch (Gemini 2.5 Flash) vs Perplexity DeepResearch
Q1 100% GR / 0% PD 60% GR / 40% SD 50% PD / 50% SD
Q2 100% GR / 0% PD 50% GR / 50% SD 80% PD / 20% SD
Q3 80% GR / 20% PD 50% GR / 50% SD 67% PD / 33% SD
Q4 75% GR / 25% PD 50% GR / 50% SD 80% PD / 20% SD
Table 5: Comparison of Deep Research Comparator with related benchmarks/platforms along key dimensions.
Benchmark/ Platform Real World User Query Long Report Fine-grained Evaluation
Deep Research Comparator
Search Arena  Miroyan et al. (2025)
Deep Research Bench   FutureSearch et al. (2025)
Deep Research Gym  Coelho et al. (2025)
BrowseComp Wei et al. (2025)
GAIA   Mialon et al. (2023)
GPQA   Rein et al. (2023)
Humanitiy’s Last Exam  HLETeam (2025)
Table 6: Scores of deep research agents on DeepResearchGym using (i) commercial search API and (ii) DeepResearchGym’s search API. Simple Deepresearch (Qwen3-8B) and Simple Deepresearch (Gemini 2.5 flash) are our baseline agent scaffold integrated with different LLMs. For GPT Researcher with commercial search API, we ran the evaluation using the newly released deep research mode, and other results are extracted from the original paper Coelho et al. (2025).
Commercial API DRGym API
System Clarity Insight Clarity Insight
Perplexity DeepResearch 89.5 89.3
gpt4-search-preview 70.1 59.1
GPT Researcher 89.8 90.4 83.7 78.0
OpenDeepSearch 59.2 47.0 61.5 49.5
HuggingFace-DeepSearch 57.5 48.0 58.3 52.4
Simple Deepresearch (Qwen3-8B) 75.6 78.1
Simple Deepresearch (Gemini 2.5 flash) 89.4 79.4

### v2

@@
 Deep Research Comparator: A Platform For Fine-grained Human Annotations of Deep Research Agents
-Abstract
-Effectively evaluating deep research agents that autonomously search the web, analyze information, and generate reports remains a major challenge, particularly when it comes to assessing long reports and giving detailed feedback on their intermediate steps.
-To address these gaps, we introduce Deep Research Comparator, a platform that offers a holistic framework for deep research agent hosting, side-by-side comparison, fine-grained human feedback collection, and ranking calculation.
-Given a user query, our platform displays the final reports from two different agents along with their intermediate steps during generation.
-Annotators can evaluate the overall quality of final reports based on side-by-side comparison, and also provide detailed feedback separately by assessing intermediate steps or specific text spans within the final report.
-Furthermore, we develop Simple Deepresearch, an end-to-end agent scaffold.
-This scaffold serves as a baseline that facilitates the easy integration of various large language models to transform them into deep research agents for evaluation.
-To demonstrate the platform’s utility for deep research agent development, we have collected real user preference data from 17 annotators on three deep research agents.
-We plan to release the code base and collected data.
-[^1] Our code is available at GitHub, and demo video can be found at Youtube.
-Deep Research Comparator: A Platform For Fine-grained Human Annotations of Deep Research Agents
-Prahaladh Chandrahasan1∗  Jiahe Jin1∗  Zhihan Zhang1∗ Tevin Wang1Andy Tang1Lucy Mo1Morteza Ziyadi2Leonardo F.R.
-Ribeiro2 Zimeng Qiu2Markus Dreyer2Akari Asai1Chenyan Xiong1 1Carnegie Mellon University  2Amazon {prahalac,jjiahe,zhihanz,tevinw,andyt,linmo}@andrew.cmu.edu {mziyadi,leonribe,zimengqi,mddreyer}@amazon.com aasai@cs.cmu.educx@cs.cmu.edu
-11footnotetext: Equal Contribution.
-1 Introduction
-Recent advancements in large language models (LLMs) have catalyzed the development of autonomous agents, which can utilize various tools to solve real-world tasks.
-Among them, the deep research agent has recently emerged, referring to an agent powered by LLMs that autonomously finds, analyzes, and synthesizes vast amounts of online information to produce a comprehensive report Xu and Peng (2025).
-Prominent proprietary examples include agents developed by OpenAI OpenAI (2025), Google Gemini Gemini (2025), and Perplexity perplexity (2025), and the open-source community has also produced notable contributions Li et al.
-(2025); GPTResearcher (2025).
-Figure 1: The main interface of Deep Research Comparator. Users can submit their query in the text box, and the intermediate steps and final reports from two agents will be displayed side by side.
-Despite these advancements, evaluating the long-form reports generated by deep research agents remains a significant challenge.
-Existing benchmarks for search agents Wei et al.
-(2025); Wu et al.
-(2025) primarily target short-form answers, leaving the assessment of long-form reports largely underexplored.
-Moreover, unlike tasks such as math or coding, there are no widely accepted objective metrics for assessing long reports (Asai et al., 2024).
-To this end, some existing work focuses on developing hand-crafted rubrics for evaluation Gou et al.
-(2025); FutureSearch et al.
-(2025); Arora et al.
-(2025), but these approaches are difficult to scale.
-Beyond the difficulty in assessing long reports, we also lack a fine-grained evaluation of agents’ intermediate steps.
-LLM-based agents often execute multi-step actions to complete long-horizon tasks, but current evaluations are typically based on outcomes Deng et al.
-(2023); Zhou et al.
-(2024); Jimenez et al.
-(2024); Xie et al.
-(2024), resulting in sparse and delayed feedback.
-However, process feedback is crucial for training methods like reinforcement learning (RL) to accurately attribute step contribution in the trajectory Feng et al.
-(2025); Wang et al.
-(2025); Setlur et al.
-(2025).
-Such feedback is also essential for understanding agent behavior and assessing the reliability of the generated reports Lightman et al.
-(2023); Wan et al.
-(2024).
-To address these gaps, we introduce Deep Research Comparator (Section 3), a platform for fine-grained human evaluation of long-report generation by deep research agents.
-It supports easy integration of new agents with minimal engineering effort, enabling both open-source and closed-source systems to be hosted through a unified web interface.
-When a user submits a query, the platform randomly routes it to two agents and presents their intermediate steps and final reports.
-Deep Research Comparator enables the collection of two types of human feedback.
-First, users can perform a side-by-side comparison of the two final reports and vote for the better one (Figure 1 bottom), similar to Chatbot Arena Chiang et al.
-(2024).
-These preference votes can then be used to calculate an outcome-based ranking of the agents.
-Second, users can provide more fine-grained annotations by upvoting or downvoting individual intermediate steps and specific text spans within the reports to indicate their quality (Figure 2).
-This detailed feedback creates a valuable dataset for potential future use, such as agent fine-tuning or reward modeling.
-To support easy deployment and evaluation of different LLMs as deep research agents, we developed Simple Deepresearch, an end-to-end agent scaffold serving as a baseline (Section 4).
-This prompt-based scaffold allows various LLMs to be easily integrated, transforming them into deep research agents for evaluation.
-Using our platform, we conducted a proof-of-concept experiment on three agents with 176 user queries (Section 5).
-We then ranked the agents based on pairwise comparison votes on their final reports.
-Additionally, we collected 1,281 upvote/downvote feedback on intermediate steps and 593 upvote/downvote with highlighted text spans in reports.
-From this, we calculated each agent’s overall upvote rate on intermediate steps for a process-based evaluation metric.
-Based on the collected data, we derive some design considerations for future agent developers on how to effectively use our platform for annotation and evaluation.
+Abstract.
+Evaluating deep research agents that iteratively search the web, analyze information, and generate reports remains a major challenge, especially in assessing long reports and providing fine-grained feedback.
+To address these gaps, we introduce Deep Research Comparator, a holistic annotation platform for the human evaluation of deep research agents.
+Our platform displays the generated reports and intermediate steps from two agents side-by-side, and allows annotators to state their preference between two final reports, and provide fine-grained feedback on specific text spans within the report or intermediate steps for each agent separately.
+Furthermore, we develop Simple Deepresearch, an agent scaffold that serves as a baseline to facilitate the integration of various large language models to transform them into deep research agents for evaluation.
+Experiment results of three agents with 17 annotators demonstrated strong alignment between human annotations and static benchmarks, while revealing divergence in fine-grained metrics, which highlights the distinct value of fine-grained annotations in providing robust evaluation and actionable data for nuanced agent optimization.
+We will provide a web-based implementation of our platform that conference attendees can interact with in real-time.
+[^1] Code and video demo are available at https://github.com/cxcscmu/Deep-Research-Comparator.
+Agentic Search, Deep Research, LLM, Human Evaluation
+copyright: nonecopyright: acmlicensedjournalyear: 2026conference: The Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Companion Proceedings of the ACM Web Conference 2026 (WWW ’26 Companion)$c$s: Computing methodologies Artificial intelligence
+1. Introduction
+Recent advancements in large language models (LLMs) have catalyzed the development of agentic search, a new paradigm in information retrieval.
+Within this paradigm, deep research agents have emerged as LLM-powered systems that autonomously find, analyze, and synthesize vast online information to produce comprehensive reports.
+Prominent commercial systems from OpenAI and Google have gained wide adoption, and the open-source community too has produced notable contributions (li2025webthinkerempoweringlargereasoning; GPTResearcher).
+Figure 1. Left: the main interface of Deep Research Comparator. Right: the fine-grained annotation mechanism for upvoting/ downvoting intermediate steps and text spans in report.
+Despite these advancements, evaluating deep research agents remains a significant challenge.
+Existing benchmarks primarily target verifiable short-form answers, leaving the assessment of long-form reports underexplored.
+While some works use hand-crafted rubrics for evaluation (gou2025mind2web2evaluatingagentic), this approach is static and hard to generalize to new tasks or domains.
+A key challenge in evaluating long report generation of deep research agents is the lack of fine-grained assessment.
+Current methods often assess reports holistically (li2025webthinkerempoweringlargereasoning; du2025deepresearch), failing to capture the uneven quality within the report, such as strong insights mixed with inaccurate content.
+This lack of text span level feedback makes it difficult to pinpoint specific areas for improvement and perform targeted agent optimization.
+Furthermore, there is a lack of process-level evaluation for the multi-step actions that agents execute.
+Current evaluations are typically based on outcomes, resulting in sparse and delayed feedback.
+However, process feedback is crucial for training methods like reinforcement learning (RL) for accurate contribution attribution.
+Such feedback is also essential for understanding agent behavior and assessing the reliability of the generated reports.
+To address these gaps, we introduce Deep Research Comparator, a platform designed to facilitate holistic human annotation for deep research agents.
+It supports easy integration of various agents with minimal engineering effort, enabling any agent system to be hosted through a unified web interface.
+The platform randomly routes user queries to two agents and presents their intermediate steps and final reports.
+Deep Research Comparator enables the collection of two types of human feedback (Figure 1).
+First, for overall report evaluation, it collects preference votes via side-by-side comparisons of final reports, similar to Chatbot Arena (chiang2024chatbot).
+Second, for fine-grained evaluation, it supports upvoting or downvoting individual intermediate steps in the generation process and specific text spans in the final reports.
+To support the evaluation of different LLMs as deep research agents, we developed Simple Deepresearch, an agent scaffold serving as a baseline, which allows various LLMs to be easily integrated to transform them into deep research agents for evaluation.
+Using our platform, we conducted a proof-of-concept experiment on three agents with 176 user queries, collecting preference votes on overall report quality and upvote/downvote feedback on 1,281 intermediate steps and 593 text spans.
+Results show human preference votes on overall report quality closely match static-benchmark evaluations, confirming the platform’s reliability.
+At the same time, the divergence in fine-grained metrics demonstrates these signals capture aspects of agent behavior that holistic scores miss.
+By providing a platform for collecting such rich fine-grained annotations, we aim to enable more robust evaluation, deeper diagnostics, and targeted optimization of deep research agents.
 Overall, we make three key contributions:
-  1. 1.
-    We provide Deep Research Comparator, an open source platform to host deep research agents and supports fine-grained human evaluation through a unified web interface.
-  2. 2.
-    We develop Simple Deepresearch, a baseline agent scaffold that allows for the easy integration of various LLMs into deep research agents for evaluation.
-  3. 3.
-    We demonstrate the utility of our platform by collecting 176 real-world user queries with comparison votes, and fine-grained annotation on 1,281 steps and 593 text spans in reports.
-2 Related Work
-Figure 2: Demonstration of fine-grained annotation.
-Left: Users can upvote/downvote each step in the generation process.
-Right: Users can highlight any text span in the final report and upvote/downvote it.
-LLM-based search agents
-Search Agents are LLM-based systems that autonomously and iteratively use search tools to gather external information Jin et al.
-(2025); Schick et al.
-(2023); Anthropic (2025); Chen et al.
-(2024) for solving complex search tasks.
-These agents can generate dynamic action decisions that are executed as real-time internet searches, and then synthesize this information to provide up-to-date and relevant responses.
-Among them, “deep research agents” represent a more capable evolution of this paradigm OpenAI (2025); GPTResearcher (2025); Li et al.
-(2025), which conduct substantially more searches, engage in long-horizon planning, and operate with significantly higher test-time compute budgets, for more demanding, research-intensive scenarios.
-Evaluations for Search Agents
-Evaluations for search agents can be roughly divided into two types: short-answer tasks and long-form report generation.
-Short-answer tasks are easier to evaluate due to the availability of definitive ground truth.
-Commonly employed evaluations include intricate tasks that demand extensive knowledge and complex reasoning, such as GPQA Rein et al.
-(2023) and Human’s Last Exam HLETeam (2025).
-Specialized benchmarks for web searches have also been developed Mialon et al.
-(2023); Wei et al.
-(2025); FutureSearch et al.
-(2025); Wu et al.
-(2025), with BrowseComp Wei et al.
-(2025) as a recent representative.
-Compared to short-answer tasks, evaluating long-form reports presents a greater challenge.
-Some efforts have introduced hand-crafted, rubric-based benchmarks Gou et al.
-(2025); Arora et al.
-(2025); FutureSearch et al.
-(2025); Asai et al.
-(2024).
-Other studies have explored using LLMs to automatically score long reports against predefined criteria Coelho et al.
-(2025); Li et al.
-(2025).
-3 Deep Research Comparator
-This section presents Deep Research Comparator, a platform for fine-grained human evaluation of diverse deep research agents.
-3.1 User Interface and Evaluation Workflow
-Our platform provides an interactive, web-based user interface designed for side-by-side, fine-grained human annotation, as shown in Figure 1.
-Annotators can submit arbitrary queries, and our system will randomly assign the query to two anonymous deep research agents.
+  1. (1)
+    We provide Deep Research Comparator, an open source platform to host deep research agents and supports fine-grained human annotation.
+  2. (2)
+    We develop Simple Deepresearch, a baseline agent scaffold that allows for the easy integration of various LLMs to transform them into deep research agents.
+  3. (3)
+    We validate Deep Research Comparator via a annotation experiment showing close alignment with static benchmarks on overall report quality and meaningful divergence in fine-grained metrics, collected from 176 real-world queries with 1,281 step and 593 span annotations.
+2. Related Work
+Search agents are LLM-based systems that autonomously use search tools to gather information (jin2025searchr1trainingllmsreason; schick2023toolformerlanguagemodelsteach; claudesearch; chen2024mindsearch) for solving complex search tasks.
+These agents can execute real-time internet searches, and synthesize information to provide responses.
+Among them, “deep research agents” represent a more capable evolution (OpenAIdeepresearch; GPTResearcher; li2025webthinkerempoweringlargereasoning), which conduct substantially more searches, engage in long-horizon planning, and operate with significantly higher test-time compute budgets, for more research-intensive scenarios.
+Evaluations for search agents fall into two types: short-answer tasks and long-form report generation.
+Short-answer tasks are easier to evaluate due to definitive ground truth.
+These include tasks demanding extensive knowledge and complex reasoning  (rein2023gpqagraduatelevelgoogleproofqa; phan2025humanitysexam), and specialized web search benchmarks (mialon2023gaiabenchmarkgeneralai; wei2025browsecompsimplechallengingbenchmark; futuresearch2025deepresearchbenchevaluating; wu2025webwalkerbenchmarkingllmsweb; wei2025browsecompsimplechallengingbenchmark).
+In contrast, evaluating long-form reports is more challenging.
+Some efforts introduced hand-crafted, rubric-based benchmarks (gou2025mind2web2evaluatingagentic; arora2025healthbenchevaluatinglargelanguage; futuresearch2025deepresearchbenchevaluating; asai2024openscholarsynthesizingscientificliterature), while other studies explored using LLMs to automatically score reports against predefined criteria (coelho2025deepresearchgymfreetransparentreproducible; li2025webthinkerempoweringlargereasoning).
+As summarized in Table 4, Deep Research Comparator is distinguished from prior work by supporting evaluation on real-world user queries, long-form reports, and fine-grained feedback mechanisms.
+3. Deep Research Comparator
+This section presents Deep Research Comparator, a platform for fine-grained human annotation of deep research agents.
+3.1. User Interface and Evaluation Workflow
+Our platform provides a web interface for side-by-side comparison and fine-grained annotation, as shown in Figure 1.
+Annotators can submit arbitrary queries, and our system will assign the query to two anonymous deep research agents.
 As the report generation begins, the agents’ intermediate steps are streamed to the front-end, followed by the final reports.
 After both reports are generated, users vote for their preferred report by selecting one of four choices: Agent A, Agent B, Tie, or Both are bad.
-These pairwise comparisons are ultimately used to compute the outcome-based rankings.
-Besides evaluating the overall quality of the final reports, this platform also supports more fine-grained evaluation (Figure 2).
-Each step of the agent is accompanied by upvote and downvote icons, allowing users to provide step-wise feedback.
-As a long report may contain content of varying quality, our platform also supports highlighting specific text spans in the final report and providing feedback either as upvotes or downvotes.
-3.2 System Architecture
-Figure 3: High-level overview of Deep Research Comparator’s system architecture.
-Figure 3 presents a high-level overview of our architecture.
+These pairwise comparisons are used to compute a ranking for overall report preference.
+Besides evaluating the overall quality of final reports, this platform also supports fine-grained annotation.
+Each step in the agents’ generation process is accompanied by upvote and downvote icons, allowing users to provide step-wise feedback.
+As a long report may contain content of varying quality, our platform also supports highlighting specific text spans in the final report and providing feedback as upvotes or downvotes.
+3.2. System Architecture
+Figure 2. High-level overview of Deep Research Comparator’s system architecture.
+Figure 2 presents an overview of our architecture.
 Our platform supports hosting a wide range of deep research agents.
-Agents that stream both intermediate steps and a final report in accordance with a unified JSON-format interface (see Appendix B) can be integrated into the platform with minimal engineering effort.
-This design allows agent developers to flexibly customize the content of intermediate steps to accommodate their annotation requirements and imposes no constraints on the agent’s internal architecture.
+Agents that generate intermediate steps and final report via a unified JSON interface can be integrated with minimal engineering effort.
+This design allows developers to flexibly customize the content of intermediate steps for their annotation requirements and imposes no constraints on the agent’s internal architecture.
 Frontend
-The frontend presents the generated intermediate steps and final reports in a consistent format on the user interface.
-It is implemented as a static web application, with static files produced during the build process and hosted in an AWS S3 bucket, served to users through CloudFront CDN caching ensuring low latency.
+The frontend presents generated intermediate steps and final reports on the user interface.
+Implemented as a static web application, its static files are produced during the build process and hosted in an AWS S3 bucket, served through CloudFront CDN caching ensuring low latency.
 Main Backend Service
 The system routes user queries from the frontend to the Main Backend Service hosted on the main CPU node.
-It is responsible for handling user requests, selecting agents, routing queries to the selected agents, processing user votes, and serving the intermediate steps, final reports, and rankings back to the frontend.
+It handles user requests, selects agents, routes queries, processes user votes, and serves intermediate steps, final reports, and rankings back to the frontend.
 Agent Serving Service
 User queries are routed to the Agent Serving Services deployed on dedicated agent CPU nodes, each encapsulated within a Docker container.
-These agent services stream their responses back to the main backend via internal API calls using the unified JSON format specified in Appendix B.
+These agent services stream their responses back to the main backend via internal API calls using the unified JSON format.
 For open-source agent services whose underlying LLMs require GPU resources, the LLMs are deployed on specialized GPU nodes and accessed by the agent services through internal API calls.
 For closed-source agents, specialized proxy pods are implemented to invoke the agents through external APIs.
-Ranking Calculation Job
-This CronJob reads agent conversations and user annotation data from the database, and computes both outcome-based evaluation metrics (such as model rankings and query counts) and process-based metrics (such as step upvote rate).
-It then writes the computed results back to the relational database, which are served to the user interface via the Main Backend Service.
-4 Simple Deepresearch
-Simple Deepresearch (Figure  4) is a prompt-based, end-to-end agent scaffold that allows for the easy integration of different LLMs via the OpenAI API, transforming it into a deep research agent.
+Metrics Calculation Job
+This CronJob reads agent conversations and user annotation data from the database, computing outcome-based evaluation metrics (like report-preference rankings) and fine-grained metrics (such as upvote rate).
+It then writes the results back to the relational database, which are served to the user interface via the Main Backend Service.
+4. Simple Deepresearch
+Simple Deepresearch is a prompt-based, end-to-end agent scaffold that allows for the easy integration of different LLMs via the OpenAI API, transforming it into a deep research agent.
 It serves as a baseline for evaluation.
-Figure 4: Overview of Simple Deepresearch, our Baseline agent scaffold.
 The agent’s workflow is an iterative process.
-At step $k$, the agent uses the user query $q$ and the history $h_{k}$ as input, and generates the thought process $t_{k}$ and action $a_{k}$.
-Each action consists of an action type and corresponding content, adhering to a predefined action space (see Table 1) specified in the agent’s system prompt (see Appendix A).
+At step $k$, the agent uses the user query $q$ and the history context $ctx_{k}$ as input, and generates the thought process $t_{k}$ and action $a_{k}$.
+Each action consists of action type and corresponding content, adhering to a predefined action space (see Table 1) specified in the agent’s system prompt.
 If $a_{k}$ is a search action, the agent queries a search API and receives documents as observation $obs_{k}$.
-Other actions do not involve interaction with the external environment but instead serve to update the agent’s memory by being incorporated into the history of next step.
-Table 1: The action space of Simple Deepresearch.
+Other actions update the agent’s memory by being incorporated into the history of next step.
+The summary action is introduced to better accommodate models with shorter context lengths.
+This allows the model to summarize the entire history, preserving the information it considers most important.
+The content of summary action will be used as history for the next step.
+Table 1. The action space of Simple Deepresearch.
 Type Content
@@
 answer The final report.
-The history $h_{k}$ contains the information of the preceding steps (from 1 to $k-1$), providing the agent with the necessary context for its current step.
-If $a_{k}$ is an answer action, then the generation process ends.
-Otherwise, the history for step $k+1$ is updated as follows:
-$h_{k+1}=\begin{cases}h_{k}+t_{k}+a_{k}+obs_{k},&a_{k}=\texttt{Search}\\ h_{k}+t_{k}+a_{k},&a_{k}=\texttt{Plan, Script}\\ a_{k},&a_{k}=\texttt{Summary}\end{cases}$
-The summary action is introduced to better accommodate models with shorter context lengths.
-This allows the model to summarize the entire history, preserving the information it considers most important.
-The action content of summary action will be used as history for the next step.
-5 Experiments
-5.1 Human Annotation
-We conduct human annotation experiments on Deep Research Comparator as proof-of-concept, with experimental setup and evaluation details provided below.
+5. Experiments
+Table 2.
+Evaluation results of deep research agents: human evaluation on Deep Research Comparator(Ranking, BT Score, Step Upvote Rate, Text Span Upvote Rate) and LLM-based evaluation on DeepResearchGym (Clarity, Insight).
+Static Benchmark Evaluation Human Annotation on Deep Research Comparator
+Agent Clarity Insight Ranking BT Score Step Upvote Rate (%) Text Span Upvote Rate (%)
+GPT Researcher 89.8 90.4 1 1135.28 88.59 74.66
+Perplexity DeepResearch 89.5 89.3 2 1087.41 90.44 90.72
+Simple Deepresearch (Gemini 2.5 Flash) 89.4 79.4 3 1000.00 78.91 90.33
+5.1. Human Annotation
+We conduct human annotation experiments as proof-of-concept, with setup and evaluation details provided below.
 Agents
-We evaluated three deep research agents: (1) the proprietary system Perplexity Deep Research perplexity (2025), (2) the open-source  GPT Researcher GPTResearcher (2025), configured in deep research mode with GPT-4.1 as the smart LLM, o4-mini as the strategic LLM, and Serper as search API (3) Simple Deepresearch(Gemini 2.5 Flash), our baseline agent scaffold intergrated with Gemini 2.5 Flash Deepmind (2025) as the underlying LLM and uses ClueWeb22 Overwijk et al.
-(2022) as search API.
+We evaluated three deep research agents: (1) the proprietary system Perplexity Deep Research, (2) the open-source  GPT Researcher (GPTResearcher), configured in deep research mode with GPT-4.1 as the smart LLM, o4-mini as the strategic LLM, and Serper as the search API (3) and Simple Deepresearch (Gemini 2.5 Flash), our baseline agent scaffold integrated with Gemini 2.5 Flash as the underlying LLM and uses ClueWeb22 as search API.
 Annotators
 A total of 17 university students and industry researchers participated in the data annotation process.
-To measure inter-annotator consistency, we also provided annotators with four fixed queries (Appendix C), which allowed us to compute human agreement on a controlled set.
+Annotators were instructed to submit their own queries to the platform that reflect significant information needs.
 Metrics
-Following Chatbot Arena Chiang et al.
-(2024), we use the Bradley-Terry model Bradley and Terry (1952) to aggregate pairwise comparison votes and compute a ranking based on user preferences for final reports.
-The final Bradley-Terry (BT) scores are computed relative to our baseline- Simple Deepresearch, which is assigned a fixed score of 1000.
-In addition to the outcome-based BT score, we also introduce a process-based evaluation metric: the step upvote rate.
-For each agent, we aggregate all upvotes and downvotes received by its intermediate steps and compute the step upvote rate as $\text{upvotes}/(\text{upvotes}+\text{downvotes})$.
-5.2 Evaluation on Static Benchmark
-To validate the performance of our baseline agent scaffold and compare human evaluation results with those from static benchmarks, we evaluated agents on DeepResearchGym Coelho et al.
-(2025), a benchmark that uses LLM-as-a-judge to assess long reports generated by deep research agents.
-Our evaluation employs two key metrics targeting different aspects of writing quality and analytical depth.
+We present two primary types of metrics computed from the platform’s annotations.
+For the evaluation of overall report quality, we follow Chatbot Arena (chiang2024chatbot) to use the Bradley-Terry (BT) model (bradley1952rank) to compute BT scores from pairwise preference votes.
+These scores are computed relative to our baseline Simple Deepresearch, which is assigned a fixed score of 1000.
+We then derive a report-preference ranking from these scores as assessment for overall report quality.
+We also introduce a fine-grained evaluation metric: the upvote rate.
+For each agent, we aggregate all upvotes and downvotes received by its intermediate steps and text spans and compute the upvote rate as $\text{upvotes}/(\text{upvotes}+\text{downvotes})$.
+It is important to note that these metrics are representative examples; Deep Research Comparator collects rich raw data (including pairwise preferences, step-wise votes, and text-span votes) that researchers can export and utilize for their own custom analyses.
+5.2. Evaluation on Static Benchmark
+To compare human evaluation for overall report quality with those from static benchmarks, we evaluated agents on DeepResearchGym (coelho2025deepresearchgymfreetransparentreproducible), a benchmark that uses LLM-as-a-judge to assess reports generated by deep research agents.
+Our evaluation employs two metrics targeting different aspects of the overall report quality.
 We provide the LLM judge with detailed evaluation criteria and prompt it to assign scores based on it.
 Specifically, Clarity assesses the logical coherence and linguistic fluency of the report, while Insight captures the analytical nuance and depth of reasoning demonstrated.
-5.3 Results
-The annotation experiment ultimately yields 176 queries across a diverse range of domains—including science, technology, media, society, economics, health, and culture—and are complex enough to require extensive web searches to answer (See Appendix D for an example).
-The win rate matrix between each pair of agents is calculated based on comparison votes on these queries, as shown in Figure 5.
-The Bradley-Terry (BT) scores and the final ranking derived from them are presented in the first two columns of Table 2.
-In addition to the side-by-side comparison votes, we collect 1,281 instances of upvote/downvote feedback on intermediate steps, and 593 instances with highlighted text spans within the reports.
-The final step upvote rate is shown in the third column of Table 2, and detailed statistics for fine-grained feedback are presented in Table 3.
-Table 2: Evaluation results of deep research agents: human evaluation on Deep Research Comparator(Ranking, BT Score, Step Upvote Rate) and LLM-based evaluation on DeepResearchGym (Clarity, Insight).
-Agent Ranking BT Score Step Upvote Rate (%) Clarity Insight
-GPT Researcher 1 1135.28 74.66 89.8 90.4
-Perplexity DeepResearch 2 1087.41 90.72 89.5 89.3
-Simple Deepresearch (Gemini 2.5 Flash) 3 1000.00 90.33 89.4 79.4
-Table 3: Upvote/downvote counts for fine-grained annotation. PD: Perplexity DeepResearch, GR: GPT Researcher, SD: Simple Deepresearch (Gemini 2.5 Flash).
-Intermediate Step Text Span
-Agent Upvote Downvote Upvote Downvote
-GR 255 33 445 151
-PR 161 17 352 36
-SD 101 27 271 29
-Figure 5: Win Probability of Agent A over Agent B.
-The results of static benchmark evaluation for agents involved in the human annotation are presented in the last two columns of Table 2, and results for additional agents are in Table 3.
-Results for human agreement are presented in Table 4.
-5.4 Discussion
-Step-Level Feedback Analysis
-Our system displays agents’ intermediate steps, with their content determined solely by agent designers.
-As a result, the content and granularity of steps vary across agents.
-Notably, although GPT Researcher ranked highest in outcome-based evaluation, it received the lowest step—upvote rate, likely due to its multi-agent parallel framework, which produces loosely ordered and often abstract or minimal step messages.
-In contrast, Simple Deepresearch (Gemini 2.5 Flash) present detailed steps that include both thought and actions (Appendix 2.
-Intermediate Steps), contributing to a higher upvote rate.
-This highlights the importance for agent designers to consider what step content to expose, tailoring the level of detail and structure to their annotation goals.
-LLM and Human Evaluation Agreement
-For report quality assessment, human rankings closely align with the LLM-based evaluations on Clarity and Insight.
-However, human agreement was relatively low.
-This may be due to the small number of queries and the fact that all three agents perform at comparably high levels (see Table 6), making fine distinctions difficult.
-This suggests two directions for annotation.
-First, as LLMs grow more capable, advancing agent performance in specialized domains may require expert annotations to provide high-quality guidance.
-Second, the subtlety of human preferences highlights the value of broader community annotations when the goal is to align agent behavior with general user expectations.
-6 Conclusion
-We present Deep Research Comparator, an open-source platform that offers a holistic framework for deep research agent hosting, web front-end demonstration, side-by-side evaluation, fine-grained annotation, and also ranking calculation.
-We also introduce Simple Deepresearch, a baseline scaffold that allows the easy integration of different LLMs to evaluate them as deep research agents.
-Using our platform, we collected votes for 176 user queries and computed rankings based on them, and also gathered fine-grained annotation for 1,281 intermediate steps and 593 text spans in reports.
-Our platform can benefit a wide range of downstream applications, such as benchmarking, agent behavior analysis, and agent training via techniques like process supervision and RLHF.
-We believe Deep Research Comparator will serve as a useful resource for the deep research agent community.
-Acknowledgements
-We would like to extend our thanks Amazon AGI for providing us with the required infrastructure support for hosting Deep Research Comparator.
+5.3. Results
+The annotation experiment yielded 176 queries across a diverse range of domains—including science, technology, society, economics, health, and culture.
+The resulting BT scores and the report-preference ranking are presented in Table 2.
+In addition to preference votes, we collected 1,281 upvote/downvote feedback instances on intermediate steps and 593 on highlighted text spans within the reports.
+The corresponding upvote rates are shown in Table 2.
+We observe a strong alignment between our human evaluation results with static benchmark evaluation.
+As shown in the first four columns, the scores from the static benchmark perfectly matches the human-derived report-preference ranking.
+However, there’s a divergence between this overall quality based ranking and the fine-grained metric upvote rates, as shown in the last four columns.
+5.4. Discussion
+The strong alignment between our human annotation and static benchmark on overall report quality validates the platform’s reliability.
+More importantly, the divergence between this overall quality and the fine-grained metrics highlights that Deep Research Comparator provides a more nuanced perspective for agent development.
+The fine-grained feedback offers actionable data beyond a single outcome score.
+For example, text-span feedback can guide fine-grained optimization of report content, while step-wise feedback provides a valuable resource for training a process reward model or data synthesis for offline reinforcement learning.
+6. Conclusion
+We present an open-source platform that offers a framework for holistic annotation for deep research agents.
+We also develop a baseline agent scaffold that allows the integration of different LLMs to evaluate them as deep research agents.
+Using our platform, we collected votes and fine-grained annotation for 176 user queries.
+We believe Deep Research Comparator will serve as a useful resource for community by benefiting a wide range of downstream applications, such as benchmarking, agent behavior analysis, and agent training via techniques like process supervision and RLHF.
+7. Ethics Statement
+All annotators were members of our research group who provided informed consent before contributing anonymous annotations, with the right to withdraw at any time.
+No personally identifiable information was collected or retained during the study.
 Appendix A Simple Deepresearch Scaffold Prompt
@@
 History Turns: (empty if this is the first turn)
-Appendix B Unified JSON format
+Appendix B History Context Update Mechanism of Simple Deepresearch
+The history context at step $k$, $ctx_{k}$, is part of the agent’s input that contains the information of the preceding steps (from 1 to k1k-1), providing the agent with the necessary information for its current step.
+If $a_{k}$ is an answer action, then the generation process ends.
+Otherwise, the history context for step $k+1$ is updated as follows:
+$ctx_{k+1}=\begin{cases}ctx_{k}+t_{k}+a_{k}+obs_{k},&a_{k}=\texttt{Search}\\ ctx_{k}+t_{k}+a_{k},&a_{k}=\texttt{Plan, Script}\\ ctx_{k},&a_{k}=\texttt{Summary}\end{cases}$
+Appendix C Unified JSON format
 # Standardized fields that all agents must support:
@@
 "citations": List[str]
-Appendix C Fixed Question Set
-  1. 1.
-    How have historical trends in Spain’s real estate market influenced the development and investment potential of golf properties?
-  2. 2.
-    How does the inclusion of personal life details in journalist biographies affect public perception of their professional credibility and integrity?
-  3. 3.
-    How does product lifecycle management affect sustainability practices within the remote car starter industry?
-  4. 4.
-    What role does digital personalization play in enhancing user engagement with their digital devices, particularly among anime fans?
 Appendix D Annotation Example
@@
 Fine-grained Annotation
-1. Text Spans
+D.1. Text Spans
 Agent: Perplexity DeepResearch
-Text Span 1 “The inclusion of personal life details in journalist biographies represents a complex intersection of transparency efforts, branding strategies, and ethical considerations in contemporary journalism.” Vote: Upvote
-Text Span 2 “Research indicates that while personal disclosures can enhance audience familiarity with reporters, they yield minimal impact on perceptions of professional credibility and integrity.” Vote: Upvote
-2. Intermediate Steps
-Agent: Simple Deepresearch (Gemini Flash 2.5)
-Intermediate Step (Part 1) Step 1 Thought: **Planning a Deep Dive into Journalists’ Credibility** Okay, so I’m tackling this interesting question: How do the personal lives of journalists impact how the public perceives their professionalism?
+Text Span 1
+“The inclusion of personal life details in journalist biographies represents a complex intersection of transparency efforts, branding strategies, and ethical considerations in contemporary journalism.” Vote: Upvote
+Text Span 2
+“Research indicates that while personal disclosures can enhance audience familiarity with reporters, they yield minimal impact on perceptions of professional credibility and integrity.” Vote: Upvote
+D.2. Intermediate Steps
+Agent: Simple Deepresearch
+Intermediate Step (Part 1)
+Step 1 Thought: **Planning a Deep Dive into Journalists’ Credibility** Okay, so I’m tackling this interesting question: How do the personal lives of journalists impact how the public perceives their professionalism?
 I’m already formulating my plan.
@@
 ACTION : Plan: 1.
-**Introduction:** * Define "journalist biographies" in this context (e.g., "About Me" pages, social media profiles, author bios in articles).
+**Introduction:** * Define ”journalist biographies” in this context (e.g., ”About Me” pages, social media profiles, author bios in articles).
 * State the core question: How does the inclusion of personal life details affect public perception of journalist credibility and integrity?
@@
 * Provide a roadmap for the article.
-Intermediate Step (Part 2) 2.
+Intermediate Step (Part 2)
+2.
 **Theoretical Framework/Literature Review:** * **Credibility:** Define and discuss established models of source credibility (e.g., expertise, trustworthiness, goodwill).
@@
 Vote: Upvote
-In our preliminary data annotation experiment, the interface only had two buttons for vote: “Agent A is better” and “Agent B is better”.
-But in the final platform we release, all four buttons are available (“Agent A is better”, “Agent B is better”, “Tie”, and “Both are bad”).
-Table 4: Human agreement results on fixed queries, measured by consensus rate.
-All submissions on our platform are anonymous, as the system does not require user login.
-Note: PD = Perplexity DeepResearch, SD = Simple Deepresearch (Gemini 2.5 Flash), GR = GPT Researcher.
-Green indicates the majority-voted agent, while red indicates the minority.
-Query GPT Researcher vs Perplexity DeepResearch Simple DeepResearch (Gemini 2.5 Flash) vs GPT Researcher Simple DeepResearch (Gemini 2.5 Flash) vs Perplexity DeepResearch
-Q1 100% GR / 0% PD 60% GR / 40% SD 50% PD / 50% SD
-Q2 100% GR / 0% PD 50% GR / 50% SD 80% PD / 20% SD
-Q3 80% GR / 20% PD 50% GR / 50% SD 67% PD / 33% SD
-Q4 75% GR / 25% PD 50% GR / 50% SD 80% PD / 20% SD
-Table 5: Comparison of Deep Research Comparator with related benchmarks/platforms along key dimensions.
+Appendix E Detailed results for Human Annotation
+Figure 3. Win Probability of Agent A over Agent B.
+Table 3. Upvote/downvote counts for fine-grained annotation. PD: Perplexity DeepResearch, GR: GPT Researcher, SD: Simple Deepresearch (Gemini 2.5 Flash).
+Intermediate Step Text Span
+Agent Upvote Downvote Upvote Downvote
+GR 255 33 445 151
+PR 161 17 352 36
+SD 101 27 271 29
+Table 4. Comparison of Deep Research Comparator with related benchmarks/platforms along key dimensions.
 Benchmark/ Platform Real World User Query Long Report Fine-grained Evaluation
 Deep Research Comparator
-Search Arena  Miroyan et al. (2025)
-Deep Research Bench   FutureSearch et al. (2025)
-Deep Research Gym  Coelho et al. (2025)
-BrowseComp Wei et al. (2025)
-GAIA   Mialon et al. (2023)
-GPQA   Rein et al. (2023)
-Humanitiy’s Last Exam  HLETeam (2025)
-Table 6: Scores of deep research agents on DeepResearchGym using (i) commercial search API and (ii) DeepResearchGym’s search API.
+Search Arena  (miroyan2025searcharenaanalyzingsearchaugmented)
+Deep Research Bench   (futuresearch2025deepresearchbenchevaluating)
+Deep Research Gym  (coelho2025deepresearchgymfreetransparentreproducible)
+BrowseComp (wei2025browsecompsimplechallengingbenchmark)
+GAIA   (mialon2023gaiabenchmarkgeneralai)
+GPQA   (rein2023gpqagraduatelevelgoogleproofqa)
+Humanitiy’s Last Exam  (phan2025humanitysexam)
+Table 5.
+Scores of deep research agents on DeepResearchGym using (i) commercial search API and (ii) DeepResearchGym’s search API.
 Simple Deepresearch (Qwen3-8B) and Simple Deepresearch (Gemini 2.5 flash) are our baseline agent scaffold integrated with different LLMs.
-For GPT Researcher with commercial search API, we ran the evaluation using the newly released deep research mode, and other results are extracted from the original paper Coelho et al.
-(2025).
+For GPT Researcher with commercial search API, we ran the evaluation using the newly released deep research mode, and other results are extracted from the original paper (coelho2025deepresearchgymfreetransparentreproducible).
 Commercial API DRGym API
@@
 Simple Deepresearch (Gemini 2.5 flash) 89.4 79.4
+Figure 4. Overview of Simple Deepresearch, our Baseline agent scaffold.